OpenStudy (anonymous):

Solve |2y - 2| = y. {2/3, 2} {2, 3} {2, 5} {1/3, 2}

hero (hero):

2y - 2 = -y 2y -2 = y

hero (hero):

2y - 2 = -y 2y + y = 2 3y = 2 y = 2/3 2y - 2 = y 2y - y = 2 y = 2 Solution: 2/3, 2